MAT 131, CALCULUS I, SUNY AT STONY BROOK, Spring 2001

Text: Calculus: concepts and contexts, J. Stewart. Syllabus: We will cover approximately Chapters 1, 2, 3, 4 and part of Chapter 5 with

certain omissions and additions of sections. A more detailed syllabus will be provided. Classdays. Classes begin on Wednesday, January 24 and end on Wednesday, May 9. Holidays: There are no classes during spring break, March 19-14. There is also no

class on Monday, April 9, due to Passover. Wednesday, April 11 will follow a Monday schedule.

Office hours: Each lecturer and teaching assistant will announce their offices hours in class.

Meetings times and locations:

Lectures

Sections Lecturer Days Time Location
1-2 Kumple MWF 4:25-5:20  Physics P118
3-4 Bishop TuTh 9:50-11:10 Harriman Hall 112
5-6 Minsky TuTh 11:20-12:40  Physics 118
7 Sung TuTh 5:30-7:20 Grad Chem 124
8 W. Kim MW 7:00-8:50  Grad Chem 126

Recitations

Sections TA Days Time Location
1  Unal TuTh 5:30-6:25 Grad Chem 126
2  Unal TuTh 3:50-4:45 Grad Chem 123
3 Leckey TuTh 11:20-12:15  Physics P129
4 Z. Han MW 3:20-4:15  Psy A 146
5 Langoni MW 11:35-12:30 Grad Chem 124
6 Z. Han MW 2:15-3:10 Grad Chem 124

Homework. A list of problems from the text will be handed out. The recitation instructors will assign, collect and return homework. The TAs may also give weekly quizzes based on the assigned problems.

Examinations. There will be two midterm tests: Tuesday, Febuary 20 and Thursday, April 5. These will be given during the regular lecture time. The final exam is Friday, Aprill 11. The location will be announced in class during the last week of the semester.

Early Midterm: There will be an evening exam at 8:30 pm on Monday, Febuary 5 in a room to be announced. This is part of the placement process and is given to all students in calculus classes. More information and a practice exam can be found at

http://www.math.sunysb.edu/earlyexam

Calculators: Calculators will not be permitted on the exams. They may be used for homework and may be useful from time to time during the semester, but are not required.

Grades. Grades are computed using: Recitation 10%, Early Midterm 10%, Midterms 20% each, Final 40%.

Disabilities. If you have any condition, such as physical or mental disability, that will make it difficult to complete the course or that will require extra time on examinations, notify your lecturer during the first two weeks of class so that appropriate arrangements can be made.

Math Learning Center: (Physics A-125/127) This is located on the A level of the physics building and provides help with pre-calculus material. The same room also contains the Calculus Resource Room where you can go for help with Calculus classes. It is staffed most days and some evenings; an exact schedule will be posted in the room. Note that this room is closed during final exam week. It also gets crowded before exams, so don’t wait until the last minute to get help there.

Calculus Online: There are many math sites on the internet. For information on the history of Calculus and short biographies of famous mathematicians, try

http://www-groups.dcs.st-and.ac.uk/~history and click on History Topics or the Bi-

ography Index. You can find information about math classes at Stony Brook at

 www.math.sunysb.edu

Drop dates: Last day for Add/Drop is Tuesday, Febuary 6. Last day to switch to lower numbered math course is Friday, March 9. The last day to withdraw from a course is Thursday, March 29.